OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3.

OpenStudy (anonymous):

please please please help

OpenStudy (amistre64):

considering the slope intercept form is: y = mx + b ; such that m=slope and b = y intercept .....

OpenStudy (amistre64):

the question i spose is designed to get you to recognize the parts of the slope intercept form

OpenStudy (anonymous):

would that be the anser cause im so confused

OpenStudy (anonymous):

ohhh hmm

OpenStudy (anonymous):

i still dont get it :(

OpenStudy (amistre64):

what dont you get?

OpenStudy (anonymous):

how would i answer this do i plug in those numbers

OpenStudy (amistre64):

"m" and "b" are a place holder, they define a generic format you are given the specific values for m and b to substitute into the general formula so yes; just replace m and b with the values stated in the problem

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

give it a shot so i can see if your understanding my directions :)

OpenStudy (anonymous):

okie lets see y=2(x)-3? or no

OpenStudy (amistre64):

perfect

OpenStudy (anonymous):

okiee enoww

OpenStudy (anonymous):

what

OpenStudy (amistre64):

usually the best course of action to take when you get to the end of a problem, is to move on to the next one ....

OpenStudy (anonymous):

oh thats it?

OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (amistre64):

parallel lines have the same slope (same m parts) that narrows it down to which two options?

OpenStudy (anonymous):

im in bit of a timed test can you give answerr please

OpenStudy (amistre64):

this is not a test answering site; this is a study site to prepare you for your tests

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

i can guide you to an answer, but you have to participate in the process

OpenStudy (amistre64):

which prolly doesnt bode well for a timed test :)